OpenStudy (anonymous):

convert r^2=4rsin tetha to cartesian

OpenStudy (anonymous):

x^2+y^2=r^2 rsin(theta)=x

OpenStudy (anonymous):

also identify the type of graph it will make....like is it a line or cirlce thx

OpenStudy (anonymous):

What would you do. Do you see?

Directrix (directrix):

x² + y² = r² sin(theta) = y/r Take this equation --> r^2 = 4r sin(theta) and substitute for r² and sin(theta). @lilshadow

OpenStudy (anonymous):

^ What he said. Same method as me.

OpenStudy (anonymous):

rsin(theta)=y sorry. Typo.

Directrix (directrix):

@Dido525 Why do you say that the "4" is not relevant?

OpenStudy (anonymous):

I mean like during the conversion.

OpenStudy (anonymous):

Of course it's still going to be part of the final answer.

OpenStudy (anonymous):

Maybe I should not have said that.

OpenStudy (anonymous):

wait guys but i get x^2+y^2=4y. What kind of graph does that make?

OpenStudy (anonymous):

Think about this in terms of Polar coordinates.

OpenStudy (anonymous):

r=4sin(theta) .

OpenStudy (anonymous):

In polar coordinates, y=sin(theta) is a circle.

OpenStudy (anonymous):

That 4 stretches it by a factor of 4.

OpenStudy (anonymous):

So in essence, you Have a circle of Radius 4.

OpenStudy (anonymous):

That's my reasoning at least.

Directrix (directrix):

@lilshadow If you don't recognize the graph from its equation, sketch the graph here: http://www.wolframalpha.com/ Enter the equation in the search field.

Directrix (directrix):

@Dido525 Isn't the diameter 4 and the radius 2?

OpenStudy (anonymous):

Yeah. Again. A silly mistake. Thanks :P .

OpenStudy (anonymous):

The radius is 2 and the Diameter is 4.

OpenStudy (anonymous):

Alrite thx guys!

Directrix (directrix):

We were happy to help.

OpenStudy (anonymous):

Yeah no problem :) .